1/2

Rabindranath Tagor,

The young prince

tradukita de Rabindranath Tagor

O mother, the young Prince is to pass by our door, – how can I attend to my work this morning? Show me how to braid up my hair; tell me what garment to put on. Why do you look at me amazed, mother? I know well he will not glance up once at my window; I know he will pass out of my sight in the twinkling of an eye; only the vanishing strain of the flute will come sobbing to me from afar. But the young Prince will pass by our door, and I will put on my best for the moment.

O mother, the young Prince did pass by our door, and the morning sun flashed from his chariot. I swept aside the veil from my face, I tore the ruby chain from my neck and flung it in his path. Why do you look at me amazed, mother? I know well he did not pick up my chain; I know it was crushed under his wheels leaving a red stain upon the dust, and no one knows what my gift was nor to whom. But the young Prince did pass by our door, and I flung the jewel from my breast before his path.

…

2/2

Traduko de la bengala poemo “Shubhokkan” de Rabindranath Tagor (∗1861-05-07 – †1941-08-07) en la Anglan de Rabindranath Tagor (∗1861-05-07 – †1941-08-07). Arg-844-1700 (2013-02-26 23:27:58) Tiu estas angligo, kiun la poeto verkis mem laŭ sia origine en la bengala lingvo verkita teksto. (Konfirmis tion al mi, Manfred Retzlaff, s-ro Probal Dasgupta.) Vidu ankaŭ la retejojn http: // eo. wikipedia. org/ wiki/ Probal_Dasgupta kaj http: // www. recmusic. org/ lieder/ get_text. html? TextId= 18070 .

http://www.poezio.net/version?poem-id=844&version-id=,1700